OpenStudy (loser66):

If \(z \in \mathbb C\), verify \(\overline {e^z}= e^{\overline {z}}\) Please, help

ganeshie8 (ganeshie8):

try using the series definition of e^z

OpenStudy (loser66):

\(e^z = \sum_{n=0}^\infty \dfrac{z^n}{n!}\) right?

OpenStudy (michele_laino):

hint: \[\Large \overline {{e^z}} = \overline {{e^{x + iy}}} = \overline {{e^x}{e^i}^y} = {e^x}{e^{ - i}}^y = ...?\]

OpenStudy (loser66):

oh, yeh!! because e^x is a real part, hence it doesn't change under the"bar" , right?

OpenStudy (loser66):

ha!! one line proof!!!

OpenStudy (michele_laino):

yes! furthermore we have the subsequent theorem: \[\Large \overline {{z_1}{z_2}} = \overline {{z_1}} \;\overline {{z_2}} \]

OpenStudy (loser66):

Thanks so much.